Solve By Substitution
y=4x+1
-8x-2y=-2

Answer by norymae(15)

substitute the y in -8x-2y=-2 by 4x+1..
-8x-2(4x+1)=-2
-8x-8x-2=-2
-16x=-2+2
-16x=0
16x/16=0/16
x=0
substitute the x in y=4x+1 by 0
y=4x+1
y=4(0)+1
y=1